# Sparse Extended Mean-Variance-CVaR Portfolios with Short-selling

### v1

@@ Abstract
This paper introduces a novel penalty decomposition algorithm customized for addressing the non-differentiable and nonconvex problem of extended mean-variance-CVaR portfolio optimization with short-selling and cardinality constraints. The proposed algorithm solves a sequence of penalty subproblems using a block coordinate descent (BCD) method while striving to fully exploit each component of the objective function or constraints. Through rigorous analysis, the well-posed nature of each subproblem of the BCD method is established, and closed-form solutions are derived where possible. A comprehensive theoretical convergence analysis is provided to confirm the efficacy of the introduced algorithm in reaching a local minimizer of this intractable optimization problem in finance, whereas generic optimization techniques either only capture a partial minimum or are not efficient. Numerical experiments conducted on real-world datasets validate the practical applicability, effectiveness, and robustness of the introduced algorithm across various criteria. Notably, the existence of closed-form solutions within the BCD subproblems prominently underscores the efficiency of our algorithm when compared to state-of-the-art methods.
@@ 1 Introduction
The mean-variance CVaR problem is crucial in finance as it provides a robust framework for optimizing portfolios by considering both expected returns (mean) and associated risks (variance and conditional value at risk). This approach helps investors balance the trade-off between maximizing returns and minimizing losses, enhancing portfolio performance and risk management in volatile markets [12, 18, 23] . Including short-selling is essential as it allows investors to profit from downward market movements, improving portfolio diversification and risk management by providing additional opportunities for profit generation in declining markets [6, 19, 24].
Moreover, sparse optimization plays a critical role in various fields such as machine learning, signal processing, and finance, offering efficient solutions for problems where the underlying data or parameters exhibit sparsity [3, 10, 13, 16, 21, 25, 26]. In portfolio optimization, having only a few assets can significantly impact diversification, potentially increasing exposure to individual asset risk and volatility, thus highlighting the importance of careful asset selection and risk management strategies. As such the cardinality constraints have been imposed in portfolio optimization models [2, 9, 14, 15, 17].
In this paper, we tackle the following extended mean-variance-CVaR portfolio problem that includes short-selling and nonconvex cardinality constraints:
$\min_{x\in\mathbb{R}^{n},\gamma\in\mathbb{R}}$ $\ f(x):=\lambda_{1}x^{T}Ax-\lambda_{2}(\Psi^{T}x-\delta\|x-\phi\|_{1})+\lambda_{3}(\gamma+\frac{1}{m(1-\beta)}\sum_{j=1}^{m}(-d_{j}^{T}x-\gamma)^{+})$ ($P$)
@@
where $\lambda_{1},\lambda_{2},\lambda_{3}=1-\lambda_{1}-\lambda_{2},\beta\in(0,1),\delta\geq 0,\eta,L,U\in\mathbb{R}^{n}$ with $L<U$ element-wise, $d_{j}\in\mathbb{R}^{n}$ for each $j\in[m],A\succeq 0,k\in[n]$. Also, we have
$\Psi=\mu+r_{c}h,\quad\eta=\mu,\quad\mbox{and}\quad d_{j}:=b_{j}-r_{c}h;\ \forall j\in[m],$
where $r_{c}\geq 0$ is the risk-neutral interest rate, $\phi_{i}$ is the proportions of the weight invested in $i$th stock and $\mu$ is the return vector. Problem ($P$) has been the subject of several research [7, 22, 20]. When $\lambda_{1}=0$, it reduces to the widely studied mean-CVaR with cardinality constraint, see for example [4, 8] and references therein.
@@
Most recently, the authors in [7] applied the penalty decomposition method (PDM) to solve ($P$) and reported promising numerical results. Nevertheless, the full structure of the problem was not exploited. Therefore, we here introduce a novel penalty decomposition algorithm customized to fully leverage every component of the objective function in ($P$) along with its constraints. We go through a sequence of penalty subproblems that a saddle point of them is efficiently identified using a BCD method. We meticulously examine each corresponding subproblem within this BCD method to derive closed-form solutions whenever feasible. Specifically, to handle the .$\|.\|_{1}$ norm in the objective function along with bound constraints, we introduce a new variable. Additionally, another variable is required to manage the softmax function in the objective function. Lastly, we resort to a third variable to address the cardinality constraint along with the short-selling constraints. Our discussions show how these new variables lead to structured subproblems and efficient We rigorously establish the convergence of our introduced algorithm towards a local minimizer of the non-differentiable problem ($P$) with non-convex constraints.
@@
Section 2 introduces Algorithm 2, a penalty decomposition algorithm designed to tackle the non-differentiable and nonconvex problem ($P$). This algorithm employs a block coordinate Algorithm 1 to address each penalty subproblem. We establish the well-posedness of these subproblems and derive closed-form solutions for three of them. The convergence analysis of both Algorithm 1 and Algorithm 2 is provided in Section 3, demonstrating that our introduced algorithm effectively reaches a local minimizer of ($P$). Additionally, Section 4 presents extensive numerical results obtained from real-world data.
@@
Notation. The complement of a set $S$ is denoted as $S^{c}$. We use $|S|$ to represent its cardinality. For a natural number $n$, we define $[n]$ as the set $\{1,2,\dots,n\}$. Now, consider a set $S$ given by $S=\{i_{1},i_{2},\dots,i_{|S|}\}$, which is a subset of $[n]$. For any vector $x$ in $\mathbb{R}^{n}$, we denote the coordinate projection of $x$ with respect to the indices in $S$ as $[x_{S};0]$, which means that the $i$th element of this vector equals to $x_{i}$ when $i$ belongs to $S$, and it equals to 00 for $i$ in $S^{c}$. We determine whether a matrix $A$ is positive semidefinite or definite by the notations $A\succeq 0$ and $A\succ 0$, respectively. In this paper, we let ${\rm sgn}(a):=1$ for $a>0,{\rm sgn}(a):=-1$ for $a<0,$ and ${\rm sgn}(0):=0$. For $x$ and $y\in\mathbb{R}^{n}$, $x\circ y$ shows the Hadamard (element-wise) multiplication of $x$ and $y$. Recall that
$\partial(\|\cdot\|_{1})|_{x}\,=\,J_{1}\times J_{2}\times\cdots\times J_{n},\quad\mbox{with}\quad J_{k}=\left\{\begin{array}[]{ll}[-1,1]&\mbox{if}\quad x_{i}=0,\\ \{{\rm sgn}(x)\}&\mbox{if}\quad x_{i}\neq 0.\\ \end{array}\right.$
@@
$\partial(\cdot)^{+}|_{x}\,=\,J_{1}\times J_{2}\times\cdots\times J_{n},\quad\mbox{with}\quad J_{k}=\left\{\begin{array}[]{ll}[0,1]&\mbox{if}\quad x_{i}=0,\\ \{1\}&\mbox{if}\quad x_{i}>0,\\ \{0\}&\mbox{if}\quad x_{i}<0.\\ \end{array}\right.$
@@ 2.1 Methodology
In this subsection, we elaborate on how we design our customized penalty decomposition algorithm. Observe that can equivalently reformulate this nonconvex problem as follows:
@@
Specifically, we introduce $y$ to deal with sparsity, $z$ to handle .$\|.\|_{1}$, and $w$ to effectively manage the last term of the objective function in ($P$). The constraints are also accordingly decoupled using these new variables such that we possibly can obtain closed-form solutions or if not, much simpler penalty subproblems in our customized penalty decomposition algorithm (see Subsection 2.2).
@@
3:  $x_{l+1}=\mbox{Argmin}_{x\in\mathcal{X}}\ q_{\rho}(x,y_{l},z_{l},w_{l},\gamma_{l})$.
4:  $y_{l+1}\in\mbox{Argmin}_{y\in\mathcal{Y}}\ q_{\rho}(x_{l+1},y,z_{l},w_{l},\gamma_{l}).$
5:  $z_{l+1}=\mbox{Argmin}_{z\in\mathcal{Z}}\ q_{\rho}(x_{l+1},y_{l+1},z,w_{l},\gamma_{l}).$
6:  $(w_{l+1},\gamma_{l+1})\in\mbox{Argmin}_{(w,\gamma)\in\mathcal{W}\times\mathbb{R}}\ q_{\rho}(x_{l+1},y_{l+1},z_{l+1},w,\gamma).$
7:  $l\leftarrow l+1$ and go to step (3).
We are now prepared to introduce our penalty decomposition algorithm, which begins with a positive penalty parameter and gradually increases it until convergence is achieved. Algorithm 1 handles the corresponding subproblem for a fixed $\rho$. For the problem ($P$), we assume having a feasible point denoted as $x^{\text{feas}}$ in hand, which is easy to obtain. To present this algorithm and its subsequent analysis, we define:
$\Upsilon\geq\max\{f(x^{\text{feas}}),\min_{x\in\mathcal{X}}q_{\rho^{(0)}}(x,y^{(0)}_{0},z^{(0)}_{0},w^{(0)}_{0},\gamma^{(0)}_{0})\}>0\qquad\text{and}\qquad X_{\Upsilon}:=\{x\in\mathbb{R}^{n}\,|\,f(x)\leq\Upsilon\}.$ (3)
@@
6:        Solve $x^{(j)}_{l+1}\in\mbox{Argmin}_{x\in\mathcal{X}}\ q_{\rho}(x,y^{(j)}_{l},z^{(j)}_{l},w^{(j)}_{l},\gamma^{(j)}_{l})$ [see (8)].
7:        Solve $y^{(j)}_{l+1}=\mbox{Argmin}_{y\in\mathcal{Y}}\ q_{\rho}(x^{(j)}_{l+1},y,z^{(j)}_{l},w^{(j)}_{l},\gamma^{(j)}_{l})$ [see (9)].
8:        Solve $z^{(j)}_{l+1}=\mbox{Argmin}_{z\in\mathcal{Z}}\ q_{\rho}(x^{(j)}_{l+1},y^{(j)}_{l+1},z,w^{(j)}_{l},\gamma^{(j)}_{l})$ [see (11)].
9:        Solve $(w^{(j)}_{l+1},\gamma^{(j)}_{l+1})\in\mbox{Argmin}_{(w,\gamma)\in\mathcal{W}\times\mathbb{R}}\ q_{\rho}(x^{(j)}_{l+1},y^{(j)}_{l+1},z^{(j)}_{l+1},w,\gamma)$ [see (16)].
@@ 2.2.1 Subproblem of $x$
Since $A\succeq 0$ and $\rho>0$, we have $A+3\rho I\succ 0$. Hence, this strictly convex problem has a unique solution, and, for $t\in\mathbb{R}$, its KKT conditions are as follows:
$12\lambda_{1}Ax_{*}-\lambda_{2}\Psi+2\rho(3x_{*}-y-z-w)+t\textbf{e}=0,\quad\mbox{and}\quad\textbf{e}^{T}x_{*}=$ 1 . (6)
Thus, $2(\lambda_{1}A+3\rho I)x_{*}=2\rho(y+z+w)+\lambda_{2}\Psi-t\textbf{e}$. By letting
$B:=\lambda_{1}A+3\rho I\qquad\text{and}\qquad\vartheta:=2\rho(y+z+w)+\lambda_{2}\Psi,$ (7)
we see $x_{*}=0.5B^{-1}(\vartheta-t\textbf{e})$, which together with $1=\textbf{e}^{T}x_{*}=0.5\textbf{e}^{T}B^{-1}(\vartheta-t\textbf{e})$ implies
$t=\frac{-1+0.5\textbf{e}^{T}B^{-1}\vartheta}{0.5\textbf{e}^{T}B^{-1}\textbf{e}},$
This in turn yields
$x_{*}=\frac{1}{2}B^{-1}\left(\vartheta+\frac{1-0.5\textbf{e}^{T}B^{-1}\vartheta}{0.5\textbf{e}^{T}B^{-1}\textbf{e}}\textbf{e}\right).$ (8)
@@ 2.2.2 Subproblem of $y$
Let $x$ and $\mu\in\mathbb{R}^{n}$ and a natural number $k\in[n]$ be given. Denote $x^{+}=\max(x,0),\,x^{-}=\max(-x,0),\,\Gamma^{+}=\{i\,|\,\eta_{i}>0\},\,\Gamma^{-}=\{i\,|\,\eta_{i}<0\},$ and $\Gamma^{0}=[n]-(\Gamma^{+}\cup\Gamma^{-})=\{i\,|\,\eta_{i}=0\}$. Then, the generalized sparsifying operator $\mathcal{S}_{k,\eta}(x)$ is defined as follows:
@@
where $\mathcal{J}$ is an index set corresponding to the $k$ largest components of $[x^{+}_{\Gamma^{+}};0]-[x^{-}_{\Gamma^{-}};0]+[x_{\Gamma^{0}};0]$ in absolute value.
@@
Proof.
For any $y$ with $\|y\|_{0}\leq k$, we can write $y=[y_{\mathcal{I}};y_{\mathcal{I}^{c}}]$ such that $y_{\mathcal{I}}=0$ for some index set $\mathcal{I}\subseteq[n]$ with $|\mathcal{I}|=n-k$. Hence, for any index set $\mathcal{I}$ with $|\mathcal{I}|=n-k$, $(P_{y})$ can be written as the following indexed problem:
$\min_{y\in\mathbb{R}^{n}}(\|x_{\mathcal{I}}-y_{\mathcal{I}}\|^{2}_{2}+\|x_{\mathcal{I}^{c}}-y_{\mathcal{I}^{c}}\|^{2}_{2})\qquad\text{subject to}\qquad{\rm sgn}(\eta_{i})y_{i}\geq 0;\quad\forall i,\quad\text{and}\quad y_{\mathcal{I}}=0,$
which is equivalent to solving:
$\min_{\theta\in\mathbb{R}^{|\mathcal{I}^{c}|}}\|\theta-x_{\mathcal{I}^{c}}\|^{2}_{2}\qquad\,\ \text{subject to}\,\qquad{\rm sgn}(\eta_{i})\theta_{i}\geq 0;\quad\forall i\in\mathcal{I}^{c}.$
For any $i\in\Gamma^{0}\cap\mathcal{I}^{c}$, we do not have any constraints, which shows that $(\theta_{*})_{i}=x_{i}$ for all $i\in\Gamma^{0}\cap\mathcal{I}^{c}$. Thus, by letting $\Gamma=\mbox{supp}(\mu)=\Gamma^{+}\cup\Gamma^{-}$, it suffices to only focus on the following:
$\min_{\nu\in\mathbb{R}^{|\Gamma\cap\mathcal{I}^{c}|}}\|\nu-x_{\Gamma\cap\mathcal{I}^{c}}\|^{2}_{2}\qquad\,\ \text{subject to}\,\qquad{\rm sgn}(\eta_{i})\nu_{i}\geq 0;\quad\forall i\in\Gamma\cap\mathcal{I}^{c}.$
The KKT conditions for the global minimizer $\nu_{*}$ of this strictly convex problem can be written as:
$x_{\Gamma\cap\mathcal{I}^{c}}=\nu_{*}-\zeta\circ({\rm sgn}(\eta))_{\Gamma\cap\mathcal{I}^{c}}\quad\text{and}\quad\nu_{*}\circ({\rm sgn}(\eta))_{\Gamma\cap\mathcal{I}^{c}}\geq 0,\quad\text{with}\quad\zeta\geq 0\quad\text{and}\quad\zeta^{T}({\rm sgn}(\eta)\circ\nu_{*})=0.$
Let $a=({\rm sgn}(\eta)\circ x)_{\Gamma\cap\mathcal{I}^{c}}$, then it is known that $a=a^{+}-a^{-}$ for $a^{+}=\max(a,0)$ and $a^{-}=\max(-a,0)$ and $a^{+}\perp a^{-}$. It is easy to see that $(\nu_{*})_{i}={a_{i}^{+}}/{{\rm sgn}(\eta_{i})}$ and $\zeta_{i}={a^{-}_{i}}$ for any $i\in\Gamma\cap\mathcal{I}^{c}$, that is,
$\nu_{*}={\rm sgn}(\eta)\circ({\rm sgn}(\eta)\circ x)^{+}_{\Gamma\cap\mathcal{I}^{c}}\qquad\text{and}\qquad\zeta=({\rm sgn}(\eta)\circ x)^{-}_{\Gamma\cap\mathcal{I}^{c}}$
satisfy the given KKT conditions above. By recalling $\Gamma^{+}$ and $\Gamma^{-}$ defined inside ($P$), we see
$(\nu_{*})_{\Gamma^{+}\cap\mathcal{I}^{c}}=x^{+}_{\Gamma^{+}\cap\mathcal{I}^{c}}\qquad\text{and}\qquad(\nu_{*})_{\Gamma^{-}\cap\mathcal{I}^{c}}=-x^{-}_{\Gamma^{-}\cap\mathcal{I}^{c}}.$
Therefore, we have
$(\theta_{*})_{\Gamma^{0}\cap\mathcal{I}^{c}}=x_{\Gamma^{0}\cap\mathcal{I}^{c}},\qquad(\theta_{*})_{\Gamma^{+}\cap\mathcal{I}^{c}}=x^{+}_{\Gamma^{+}\cap\mathcal{I}^{c}},\qquad\text{and}\qquad(\theta_{*})_{\Gamma^{-}\cap\mathcal{I}^{c}}=-x^{-}_{\Gamma^{-}\cap\mathcal{I}^{c}}.$
So, for any index $\mathcal{I}$ specified above, the optimal value of the indexed problem is given by
$\|x_{\mathcal{I}^{c}}-\theta_{*}\|^{2}_{2}+\|x_{\mathcal{I}}\|^{2}_{2}$ $=$ $\|x_{\Gamma^{+}\cap\mathcal{I}^{c}}-x^{+}_{\Gamma^{+}\cap\mathcal{I}^{c}}\|^{2}_{2}+\|x_{\Gamma^{+}\cap\mathcal{I}}\|^{2}_{2}$
$+$ $\|x_{\Gamma^{-}\cap\mathcal{I}^{c}}+x^{-}_{\Gamma^{-}\cap\mathcal{I}^{c}}\|^{2}_{2}+\|x_{\Gamma^{-}\cap\mathcal{I}}\|^{2}_{2}$
$+$ $\|x_{\Gamma^{0}\cap\mathcal{I}^{c}}-x_{\Gamma^{0}\cap\mathcal{I}^{c}}\|^{2}_{2}+\|x_{\Gamma^{0}\cap\mathcal{I}}\|^{2}_{2},$
where
$\|x_{\Gamma^{+}\cap\mathcal{I}^{c}}-x^{+}_{\Gamma^{+}\cap\mathcal{I}^{c}}\|^{2}_{2}+\|x_{\Gamma^{+}\cap\mathcal{I}}\|_{2}^{2}=\|x_{\Gamma^{+}\cap\mathcal{I}^{c}}\|_{2}^{2}-\|x^{+}_{\Gamma^{+}\cap\mathcal{I}^{c}}\|^{2}_{2}+\|x_{\Gamma^{+}\cap\mathcal{I}}\|_{2}^{2}=\|x_{\Gamma^{+}}\|^{2}_{2}-\|x^{+}_{\Gamma^{+}\cap\mathcal{I}^{c}}\|_{2}^{2},$
$\|x_{\Gamma^{-}\cap\mathcal{I}^{c}}+x^{-}_{\Gamma^{-}\cap\mathcal{I}^{c}}\|^{2}_{2}+\|x_{\Gamma^{-}\cap\mathcal{I}}\|_{2}^{2}=\|x_{\Gamma^{-}\cap\mathcal{I}^{c}}\|_{2}^{2}-\|x^{-}_{\Gamma^{-}\cap\mathcal{I}^{c}}\|^{2}_{2}+\|x_{\Gamma^{-}\cap\mathcal{I}}\|_{2}^{2}=\|x_{\Gamma^{-}}\|^{2}_{2}-\|x^{-}_{\Gamma^{-}\cap\mathcal{I}^{c}}\|_{2}^{2},$
and
$\|x_{\Gamma^{0}\cap\mathcal{I}^{c}}-x_{\Gamma^{0}\cap\mathcal{I}^{c}}\|^{2}_{2}+\|x_{\Gamma^{0}\cap\mathcal{I}}\|_{2}^{2}=\|x_{\Gamma^{0}}\|^{2}_{2}-\|x_{\Gamma^{0}\cap\mathcal{I}^{c}}\|_{2}^{2}.$
Consequently, the optimal value becomes
$\|x_{\mathcal{I}^{c}}-\theta_{*}\|^{2}_{2}+\|x_{\mathcal{I}}\|^{2}_{2}$ $=$ $\|x_{\Gamma^{+}}\|^{2}_{2}-\|x^{+}_{\Gamma^{+}\cap\mathcal{I}^{c}}\|_{2}^{2}$
$+$ $\|x_{\Gamma^{-}}\|^{2}_{2}-\|x^{-}_{\Gamma^{-}\cap\mathcal{I}^{c}}\|_{2}^{2}$
$+$ $\|x_{\Gamma^{0}}\|^{2}_{2}-\|x_{\Gamma^{0}\cap\mathcal{I}^{c}}\|_{2}^{2}$
$=$ $\|x\|^{2}_{2}-\|x^{+}_{\Gamma^{+}\cap\mathcal{I}^{c}}\|_{2}^{2}-\|x^{-}_{\Gamma^{-}\cap\mathcal{I}^{c}}\|_{2}^{2}-\|x_{\Gamma^{0}\cap\mathcal{I}^{c}}\|_{2}^{2}.$
[…]

### v2

@@ Abstract
This paper introduces a novel penalty decomposition algorithm customized for addressing the non-differentiable and nonconvex problem of extended mean-variance-CVaR portfolio optimization with short-selling and cardinality constraints. The proposed algorithm solves a sequence of penalty subproblems using a block coordinate descent (BCD) method while striving to fully exploit each component within the objective function and constraints. Through rigorous analysis, the well-posedness of each subproblem of the BCD method is established, and closed-form solutions are derived where possible. A comprehensive theoretical convergence analysis is provided to confirm the efficacy of the introduced algorithm in reaching a Lu–Zhang minimizer for this intractable optimization problem. Numerical experiments conducted on real-world datasets validate the practical applicability and effectiveness of the introduced algorithm based on various criteria. Notably, the existence of closed-form solutions within the BCD subproblems prominently underscores the efficiency of our algorithm when compared to state-of-the-art methods.
@@ 1 Introduction
The mean-variance CVaR problem is crucial in finance, as it provides a robust framework for optimizing portfolios by considering both expected returns (mean) and associated risks (variance and conditional value at risk). This approach helps investors balance the trade-off between maximizing returns and minimizing losses, enhancing portfolio performance and risk management in volatile markets [5, 13, 18]. Including short-selling is essential, as it allows investors to profit from downward market movements, improving portfolio diversification and risk management by providing additional opportunities for profit generation in declining markets [20].
Moreover, sparse optimization plays a critical role in various fields, such as machine learning, signal processing, and finance, offering efficient solutions for problems where the underlying data or parameters exhibit sparsity [3, 6, 19, 22, 21]. In portfolio optimization, having only a few assets can significantly impact diversification, potentially increasing exposure to individual asset risk and volatility. This highlights the importance of careful asset selection and risk management strategies. Hence, the cardinality constraints have been imposed in various portfolio optimization models [4, 11, 14].
In this paper, we tackle the following extended mean-variance-CVaR portfolio problem, which includes short-selling and nonconvex cardinality constraints:
$\min_{x\in\mathbb{R}^{n},\gamma\in\mathbb{R}}$ $\ f(x,\gamma)=\lambda_{1}x^{T}Ax-\lambda_{2}(\Psi^{T}x-\delta\|x-\phi\|_{1})+\lambda_{3}(\gamma+\frac{1}{m(1-\beta)}\sum_{j=1}^{m}(-d_{j}^{T}x-\gamma)^{+})$ ($P$)
@@
where $\lambda_{1}$ and $\lambda_{2}\geq 0$, and $\lambda_{3}=1-\lambda_{1}-\lambda_{2}\geq 0$ are the weights of variance, return, and CVaR, respectively. $\beta\in(0,1)$ is the confidence level, $\delta\geq 0$ is the cost per change in the proportion of the $i$-th asset, $L,U\in\mathbb{R}^{n}$ and $L<U$ are the lower and upper bounds for the weight vector of assets $x$, and $A\succeq 0$ is the covariance matrix. Also we have $\Psi=\mu+r_{c}h,\eta=\mu$, and $d_{j}:=b_{j}-r_{c}h;\ \forall j\in[m]$, where $\mu$ is the return vector, $r_{c}\geq 0$ is the risk-neutral interest, $h_{j}$ is the portion of risk free return for the investor when stock $i$ is on a short selling position. Finally, $b_{j},\ j\in[m]$ are the generated scenarios and $\phi_{i}$ is the proportion of the initial wealth invested in the $i$th stock.
Problem ($P$) has been the subject of several research [9, 16, 17]. When $\lambda_{1}=0$, it reduces to the widely studied mean-CVaR with a cardinality constraint, see, for example, [7, 10] and references therein. It is worth noting that by minimizing both Variance and CVaR simultaneously, we are essentially seeking a portfolio that exhibits low overall volatility (variance) and resilience to crises (CVaR). In the event of a major market shock (such as the 2020 COVID crash), the portfolio is explicitly designed to limit the average severity of losses. This is why we use two measures of risk together.
@@
Recently, Hamdi et al. [9] applied the penalty decomposition method (PDM) to solve ($P$) and reported promising numerical results. Nevertheless, the full structure of the problem was not exploited. Therefore, we introduce a novel penalty decomposition algorithm customized to fully leverage every component of the objective function in ($P$) along with its constraints. We solve a sequence of penalty subproblems, of which a saddle point is efficiently identified using a BCD method. We meticulously examine each corresponding subproblem within this BCD method to derive closed-form solutions whenever feasible. Specifically, to handle the $\|.\|_{1}$ norm in the objective function along with bound constraints, we introduce a new variable. Additionally, another variable is required to manage the softmax function in the objective function. Lastly, we resort to a third variable to address the cardinality constraint, along with the short-selling constraints. Our discussions show how introducing these new variables leads to structured subproblems, which, in turn, admit closed-form or otherwise efficient solutions. We rigorously establish the convergence of our introduced algorithm toward a Lu–Zhang [12] minimizer of the non-differentiable problem ($P$) with non-convex constraints.
@@
Section 2 introduces Algorithm 2, a penalty decomposition algorithm designed to tackle the non-differentiable and nonconvex problem ($P$). This algorithm employs a block coordinate Algorithm 1 to address each penalty subproblem. We establish the well-posedness of these subproblems and derive closed-form solutions for three of them. The convergence analysis of both Algorithm 1 and Algorithm 2 is established in Section 3, demonstrating that our introduced algorithm effectively reaches a Lu–Zhang minimizer of ($P$). Additionally, Section 4 presents extensive numerical results obtained from real-world data. To improve the readability of the paper, we have moved the proofs of some lemmas and theorems to the appendix.
@@
Notation. The complement of a set $S$ is denoted as $S^{c}$. We use $|S|$ to represent its cardinality. For a natural number $n$, we define $[n]$ as the set $\{1,2,\dots,n\}$. Now, consider a set $S$ given by $S=\{i_{1},i_{2},\dots,i_{|S|}\}$, which is a subset of $[n]$. For any vector $x$ in $\mathbb{R}^{n}$, we denote the coordinate projection of $x$ with respect to the indices in $S$ as $x_{S};0]$, which means that the $i$th element of this vector equals $x_{i}$ when $i$ belongs to $S$, and it equals 0 for $i$ in $S^{c}$. We determine whether a matrix $A$ is positive semidefinite or definite by the notations $A\succeq 0$ and $A\succ 0$, respectively. In this paper, we let ${\rm sgn}(a):=1$ for $a>0,{\rm sgn}(a):=-1$ for $a<0,$ and ${\rm sgn}(0):=0$. For $x$ and $y\in\mathbb{R}^{n}$, $x\circ y$ shows the Hadamard (element-wise) multiplication of $x$ and $y$. Recall that [2]
[…]
@@
[…]
@@ 2.1 Methodology
In this subsection, we elaborate on how we design our customized penalty decomposition algorithm. Observe that we can equivalently reformulate this nonconvex problem as follows:
@@
Specifically, we introduce $y$ to deal with sparsity, $z$ to handle $\|.\|_{1}$, and $w$ to effectively manage the last term of the objective function in ($P$). The constraints are also decoupled accordingly using these new variables, so that we can possibly obtain closed-form solutions or, if not, much simpler penalty subproblems in our customized penalty decomposition algorithm (see Subsection 2.2).
@@
[…]
We are now prepared to introduce our penalty decomposition algorithm, which begins with a positive penalty parameter and gradually increases it until convergence is achieved. Algorithm 1 handles the corresponding subproblem for a fixed $\rho$. For the problem ($P$), we assume that we have a feasible point denoted as $x^{\text{feas}}$ in hand, which is easy to obtain. To present this algorithm and its subsequent analysis, we define:
$\Upsilon$ $\geq\;\max\!\Big\{f(x^{\text{feas}},\gamma^{\text{feas}}),\;\min_{x\in\mathcal{X}}q_{\rho^{(0)}}\!\big(x,y^{(0)}_{0},z^{(0)}_{0},w^{(0)}_{0},\gamma^{(0)}_{0}\big)\Big\}\;>0,$
$X_{\Upsilon}$ $=\;\Big\{(x,\gamma)\in\mathbb{R}^{n}\times\mathbb{R}\ \Big|\ f(x,\gamma)\leq\Upsilon\Big\}.$
@@
6:   Solve $x^{(j)}_{l+1}=\mbox{Argmin}_{x\in\mathcal{X}}\ q_{\rho^{(j)}}(x,y^{(j)}_{l},z^{(j)}_{l},w^{(j)}_{l},\gamma^{(j)}_{l})$ [see (5)].
7:   Solve $y^{(j)}_{l+1}\in\mbox{Argmin}_{y\in\mathcal{Y}}\ q_{\rho^{(j)}}(x^{(j)}_{l+1},y,z^{(j)}_{l},w^{(j)}_{l},\gamma^{(j)}_{l})$ [see (7)].
8:   Solve $z^{(j)}_{l+1}=\mbox{Argmin}_{z\in\mathcal{Z}}\ q_{\rho^{(j)}}(x^{(j)}_{l+1},y^{(j)}_{l+1},z,w^{(j)}_{l},\gamma^{(j)}_{l})$ [see (8)].
9:   Solve $(w^{(j)}_{l+1},\gamma^{(j)}_{l+1})\in\mbox{Argmin}_{(w,\gamma)\in\mathcal{W}\times\mathbb{R}}\ q_{\rho^{(j)}}(x^{(j)}_{l+1},y^{(j)}_{l+1},z^{(j)}_{l+1},w,\gamma)$ [see (9)].
@@ 2.2.1 Subproblem of $x$
Lemma 2.1.
The solution to the problem ($P_{x}$) is
$x_{*}=\frac{1}{2}B^{-1}\left(\vartheta+\frac{1-0.5\textbf{e}^{T}B^{-1}\vartheta}{0.5\textbf{e}^{T}B^{-1}\textbf{e}}\textbf{e}\right),$ (5)
where
$B:=\lambda_{1}A+3\rho I\qquad\text{and}\qquad\vartheta:=2\rho(y+z+w)+\lambda_{2}\Psi.$ (6)
@@ 2.2.2 Subproblem of $y$
Let $x$ and $\eta\in\mathbb{R}^{n}$ and a natural number $k\in[n]$ be given. Denote $x^{+}=\max(x,0),\,x^{-}=\max(-x,0),\,\Gamma^{+}=\{i\,|\,\eta_{i}>0\},\,\Gamma^{-}=\{i\,|\,\eta_{i}<0\},$ and $\Gamma^{0}=[n]-(\Gamma^{+}\cup\Gamma^{-})=\{i\,|\,\eta_{i}=0\}$. Then, the generalized sparsifying operator $\mathcal{S}_{k,\eta}(x)$ is defined as follows:
@@
where $\mathcal{J}$ is an index set corresponding to the $k$ largest components of $x^{+}_{\Gamma^{+}};0]-[x^{-}_{\Gamma^{-}};0]+[x_{\Gamma^{0}};0]$ in absolute value.